A small particle has charge -2.00 μC and mass 2.60×10−4 kg . It moves from point A, where the electric potential is VA = 270 V , to point B, where the electric potential VB = 840 V is greater than the potential at point A. The electric force is the only force acting on the particle. The particle has a speed of 3.10 m/s at point A. What is its speed at point B?
Answer #1

  ________________________________________...

Charge = q = -2.00*10^-6 C

Mass = m = 2.60×10−4kg

Initial velocity = u =3.10 m/s

Final velocity = y

Change in kinetic energy = change in potential energy

(1/2)mv^2 - (1/2)mu^2 = q (Va - Vb)

multiply by 2 abd divide by m on both sides

v^2 - u^2 = 2q (Va - Vb) /m

v^2 = u^2 + 2q (Vb - Va) /m

v = sq rt [ u^2 + 2q (Va - Vb) /m ]

v = sq rt [ 3.1^2 - 2*2.00*10^-6 (270 -840) / 2.60×10−4 ]

v = sq rt [ 9.61 +4.00*10^-6 (570 ) / 2.60×10−4 ]

v = sq rt [ 9.61 +4.00*10^-2 (285 ) ]

v = sq rt [ 9.61 + 11.4 ]

v = sq rt [ 21.04 ]

v = 4.58 m/s

A) Its speed at point B is 4.58 m/s

B)It is moving faster at B than at A